A person with hypermetropia has a near-point of 35 cm and a lens-to-retina distance of 2.0 cm.

What is the maximum optical power of this person's eye? ( to 2 s.f and in D)
1
Expert's answer
2020-05-29T09:51:21-0400

Optical power is a reciprocal of the focal length. We can find the focal length according to thin lens equation:


"P=\frac{1}{f}=\frac{1}{i}+\frac{1}{o}=\frac{1}{0.2}+\frac{1}{0.35}=53D."